OpenStudy (anonymous):

How would you solve 441x⁴ - 361y²

OpenStudy (raden):

hint : 441 = 21^2 and 361=19^2 i think u need the factors not to find the solution, right ?

OpenStudy (anonymous):

Yea, but factoring confuses me -.-

OpenStudy (raden):

u might can use the property : a^2b^2 - m^2n^2 = (ab)^2 - (mn)^2 = (ab+mn)(ab-mn)

OpenStudy (anonymous):

Okay...I'll attempt

OpenStudy (raden):

dont forget, x^4 = (x^2)^2 :)

OpenStudy (anonymous):

Okay thx :)

OpenStudy (raden):

welcome..

OpenStudy (anonymous):

 How would i solve this

OpenStudy (raden):